)An Iron-Constantan thermocouple (TC) is used to measure temperature between 0 o C and 300 oC at which it generates 5268 μV. If the thermal emf is 12500 μV relative to the reference junction at 20 o C. Estimate the measured junction temperature.


1
Expert's answer
2021-01-23T10:29:19-0500

Let's solve this problem:

"0^o C \space \space to \space \space 300^oC \space \space \space \space 5268\mu V"

"20^o C \space \space to \space \space T_k \space ^oC \space \space \space \space12500\mu V"

The electro motive force : "E = K(T_k-T_C)"

Where: K is the "\frac{T}{C}" sensitivity

"5268 = K(300-0)"

"12500 = K(T_k - 20)" ------> "\large\frac{5268}{12500} = \large\frac{300}{T_k-20}"

"T_k = \large\frac{12500*300-20*5268}{5268}" "= 691.845 ^oC"